# FUSAR-KLIP: Towards Multimodal Foundation Models for Remote Sensing

### v2

@@ Abstract
Cross-modal artificial intelligence has garnered widespread attention in recent years, achieving significant progress in the study of natural images. However, existing methods are mostly designed for RGB imagery, leaving a significant gap in modeling synthetic aperture radar (SAR) imagery. SAR, with its all-day, all-weather imaging capabilities, plays an irreplaceable role in remote sensing scene understanding. To address this gap, this paper proposes FUSAR-KLIP, the first universal SAR multimodal foundational model, along with reusable data and evaluation baselines. Specifically: (1) This work introduces the critical yet long-overlooked attribute of geographic information into remote sensing research, constructing FUSAR-GEOVL-1M (the first large-scale SAR dataset with complete geographic projection properties), covering multiple satellite platforms, 120,000 images, and 135 cities. (2) Aligned structured text is generated through a hierarchical cognitive chain-of-thought (HCoT), providing more than one million multi-dimensional semantic annotations of landforms, regional functions, target attributes, and spatial relationships. (3) We design a Self-Consistent Iterative Optimization mechanism that continuously enhances cross-modal alignment through a self-supervised closed loop of contrastive, matching, and reconstruction learning on a transferable multimodal encoder. (4) A unified evaluation benchmark is established across 11 representative downstream vision and vision-language tasks, with comparisons against 14 leading foundation models, where FUSAR-KLIP demonstrates leading performance, particularly in object counting and land-cover classification. We expect that FUSAR-KLIP’s large-scale multimodal data, transferable model architecture, and comprehensive experimental benchmark will significantly advance the development of SAR multimodal baseline models. The dataset and model are publicly available at: https://github.com/yangyifremad/SARKnowLIP.
@@ 1 Introduction
Despite the remarkable progress of cross-modal artificial intelligence in natural imagery and optical remote sensing[1],[2], the applicability of existing visual–language models (VLMs) to synthetic aperture radar (SAR) imagery remains severely limited. Most prevailing approaches are built upon modeling assumptions designed for RGB imagery[3],[4],[5],[6],[7], which fail to effectively capture the unique imaging mechanisms and semantic characteristics of SAR data. As a result, direct transfer often leads to inadequate representation and performance degradation. Consequently, dedicated multimodal models for SAR are still largely absent, constituting a critical bottleneck that constrains the advancement of cross-modal remote sensing research toward higher-level semantic understanding.
This situation contrasts sharply with the strategic role of SAR in the remote sensing ecosystem. As an active imaging modality, SAR provides all-weather, day-and-night observation capabilities and exhibits heightened sensitivity to structural and geometric properties of ground objects[8], enabling reliable information acquisition under nighttime, extreme weather, and high cloud or precipitation conditions. Owing to these characteristics, SAR has been widely applied to disaster monitoring, oceanic and polar observation, urban management, and national security[9]. Compared with optical and multispectral imagery, SAR is better suited for long-term monitoring in complex environments, and its irreplaceable value highlights the urgent need for dedicated multimodal foundation models tailored to SAR[10].
However, due to limitations of SAR in terms of data privacy, domain expertise, and complex features, this research direction still faces many challenges. Specifically:
@@
1) Geospatial Priors Remain Underutilized: Geographic information, a core element in Earth sciences, is crucial for RS. However, most current SAR interpretation studies remain at the visual level, neglecting geographic attributes. This stems from the extension of natural image paradigms—focused mainly on texture and structure[11]—as well as the closed nature of SAR data sources, with many studies still relying on outdated datasets like MSTAR from the 1990s[12]. As shown in Table I, our review of 16 mainstream SAR datasets across tasks such as detection, segmentation, and recognition reveals a widespread omission of geographic metadata. This absence hinders regional analysis, terrain understanding, and spatial reasoning, ultimately limiting progress toward cognitive-level SAR interpretation.
@@
2) Knowledge Bottleneck: Existing RS VLM studies face a significant knowledge bottleneck in text construction. On the one hand, approaches that convert detection, recognition, or segmentation annotations into templated text [13],[14] generate descriptions that are semantically sparse and structurally repetitive, covering only shallow attributes (e.g., category and location) while neglecting complex elements such as landform, regional function, or spatial structure. This low-information text causes different images to collapse into similar semantic spaces, weakening alignment quality and fine-grained discrimination, as shown in Fig. 2(a). On the other hand, automatic text acquisition strategies, which are effective in natural image or optical RS domains (e.g., web crawling, video subtitles, or AI-generated captions), fail in the SAR domain due to speckle noise and the need for expert knowledge. The resulting AI-generated text is often semantically vague or distorted Fig. 2(b). Together, these issues highlight that SAR multimodal foundational model suffers from a dual bottleneck: text derived from manual annotations lacks semantic depth, and reliable large-scale automatic acquisition remains infeasible, severely restricting scalability in open-world applications.
@@
FUSAR-MAP[16] 2021 45 C Single 1 1,024 610 TIF Segmentation
SSDD[17] 2021 511 C,X Multi $\sim1$5 500 1,160 JPG Detection
@@
FUSAR-SHIP[22] 2022 271 C Single 1 512 16,144 TIFF Detection
@@
SAMPLE[24] 2019 14 X Single 0.3 128 2,732 PNG Classification
@@
PolarSAR-Seg[27] 2022 36 C Multi 8 512 500 TIFF Segmentation
@@
In terms of text data acquisition, this paper introduces a hierarchical cognitive chain-of-thought (HCoT) instruction to simulate the human interpretation process of SAR imagery. HCoT guides GPT-4.1 [30] to progressively incorporate multi-dimensional knowledge—such as geographic context, regional priors, SAR imaging principles, and target scale perception—enabling semantic reasoning from global to local, and from general to domain-specific knowledge. Furthermore, the constructed multi-scale image semantics driving mechanism guides the model to generate semantic expressions that connect the upper and lower scales on large, medium, and small scale images, and establish cognitive coherence. Under HCoT guidance, the language model significantly enhances its SAR interpretation ability, generating more informative and coherent textual descriptions to support multimodal representation learning.
In terms of multimodal alignment modeling, this paper proposes FUSAR-KLIP, a dual-tower multimodal framework employing Vision Transformer (ViT)[31] and BERT[32] as visual and language encoders. The model jointly optimizes image-text contrastive loss (ITC), matching loss (ITM), and masked language modeling loss (MLM) to construct a unified cross-modal embedding space bridging low-level perception and high-level semantics. To mitigate semantic deviations in HCoT-guided generated text, we introduce a self-consistent iterative optimization (SCIO) module that enhances alignment accuracy and stability through a closed-loop self-supervised strategy of screening, filtering, and reconstruction.
@@
  • FUSAR-KLIP(Knowledge-guided Language-Image Pre-training): The first visual language foundational model for SAR is developed, combining contrast, matching, and reconstruction multi-task learning to establish a cross-modal representation space for vision-language collaboration. The SCIO optimization module is introduced to dynamically enhance text accuracy and improve cross-modal alignment quality through the ”screen–filter–reconstruct” self-supervised closed-loop optimization mechanism.
  • Leading Multi-Task Generalization Capability: In various typical downstream tasks, such as target classification, detection, land feature segmentation, image captioning, image-text retrieval, and visual question answering, FUSAR-KLIP demonstrates superior semantic understanding and cross-task generalization performance compared to existing RS multimodal models.
The remainder of this paper is organized as follows: Section 2 introduces the research progress in related fields; Section 3 details the construction process of the FUSAR-GEOVL-1M dataset and the design of the multimodal self-supervised model; Section 4 presents experimental verification and analysis across multiple RS downstream tasks; Section 5 provides a summary of the paper and outlines potential future research directions.
@@ 2.1 Paradigm Evolution of Remote Sensing Interpretation
In the supervised stage, large-scale manual annotations enabled significant progress in recognition, detection, and segmentation tasks, but the prohibitive cost and limited semantic coverage of annotations severely constrained generalization across diverse scenes.
Unimodal self-supervised learning (SSL) alleviated this bottleneck by mining latent structures from unlabeled data, producing robust representations with minimal fine-tuning and laying the foundation for domain-specific models[33]. Nevertheless, unimodal SSL remains largely restricted to low-level perceptual cues such as texture and geometry, offering limited capacity for higher-order semantic reasoning[34],[35].
Multimodal SSL further introduces language to build cross-modal embedding spaces through large-scale image–text alignment, enabling richer semantics and task generalization. This paradigm has driven breakthroughs in natural imagery and optical remote sensing. However, SAR has received little attention within this framework: most existing models rely on RGB-based assumptions and fail to capture SAR’s unique imaging mechanisms. As a result, researchers have only begun to explore unimodal SSL in SAR as a transitional step toward multimodal expansion, leaving the development of SAR-specific multimodal foundation models an urgent and open challenge.
@@ 2.2 Research on Unimodal Self-supervised Foundational Models in SAR Images
SARATR-X uses a two-step SSL method with multi-scale gradient features to establish a high-performance SAR image target recognition foundational model[28]. Yang et al. proposed the SARDet-CL, combining feature enhancement with a SSL method constrained by imaging mechanisms, and achieved advanced performance in downstream detection tasks[36]. Li et al. proposed SAR-JEPA, which constructed a self-supervised pre-training foundational model for SAR target recognition tasks from the perspective of overcoming speckle noise[37]. Ren et al. introduced multi-image factor SSL to promote directional feature learning and obtain generalized features, enhancing the performance in terrain classification tasks[38]. Wang et al. constructed a SAR multi-task foundation model based on cross-domain continuous pre-training[39]. Pei et al. proposed a two-stage SAR image pre-training method based on SSL to improve the accuracy of target classification[40]. MSFA proposed a multi-stage filter augmentation pre-training framework for large-scale RGB and SAR data, which performed well when transferred to detection tasks[23].
@@ 2.3 Research on Multimodal VLMs in Remote Sensing Field
Existing work primarily focuses on optical imagery. For instance, RemoteCLIP is based on the CLIP framework by converting class labels in public RS datasets into templated text, establishing a foundational model that achieves leading performance across many typical tasks[13]. SkyEyeGPT developed a multimodal RS command dataset and designed a two-stage tuning strategy to enhance conversational capabilities[42]. VHM generated the HqDC dataset based on the large language model Gemini, easing the issue of model hallucination[43]. SkyScript utilized structured geographic information from OpenStreetMap to generate semantic text for optical images, improving the quality of cross-modal alignment[44]. GeoChat expanded existing RS image-text pairs to build a multi-round conversation dataset with command-following capabilities[45]. BITA introduced a lightweight Fourier Transformer structure to enhance image-text interaction[46], while BAN developed a foundational model for RS change detection tasks, leading to improved task performance[47]. Multimodal modeling for SAR images is still in its infancy.
In contrast, multimodal modeling for SAR is still in its infancy: early attempts such as SARLANG[14], which generates text from detection annotations, and SSL-LIP[48], which employs dual-stage self-supervised training with limited labels, have demonstrated feasibility but remain constrained by small-scale, template-based text construction and sparse semantics. These limitations highlight the urgent need for a scalable multimodal foundational model, grounded in self-supervised learning and tailored to SAR’s unique imaging characteristics, to advance higher-level semantic modeling and intelligent interpretation.
@@ 3.1 FUSAR-GEOVL-1M Image Data
The dataset covers 135 representative urban areas, as shown in Fig. 5. The scene types include typical RS semantic categories such as airports, ports, urban areas, water bodies, industrial parks, and road networks. The dataset construction process consists of the following key steps:
@@
As shown in Fig. 4, the FUSAR-GEOVL-1M dataset presents a clear distribution of data, sample proportions, and representative examples. It is noteworthy that this dataset is built upon several task-oriented SAR datasets previously proposed by our team—namely, FUSAR-MAP[16], FUSAR-SHIP[22], and FUSAR-AIR[54]—which have undergone large-scale expansion and data restructuring. These enhancements have significantly improved the dataset’s coverage, semantic density, and adaptability to downstream tasks. The aforementioned datasets have been downloaded over 6,000 times and are widely cited and applied in the field of SAR interpretation. FUSAR-GEOVL-1M represents the systematic accumulation of our team’s long-term research in SAR image understanding and provides a strong data foundation for multimodal modeling in the SAR domain.

### v3

@@ Abstract
Cross-modal artificial intelligence, represented by visual language models, has achieved significant success in general image understanding. However, a fundamental cognitive inconsistency exists between general visual representation and remote sensing image interpretation: remote sensing images couple topography, terrain, and spatial structure, thereby inherently requiring models to possess deep geoscientific understanding. This cognitive difference is further amplified in synthetic aperture radar (SAR) imagery: while SAR possesses irreplaceable all-weather, all-day observation capabilities, it is constrained by coherent imaging mechanisms, exhibiting significant modal heterogeneity with general images. To address this inconsistency, we propose FUSAR-KLIP, the first knowledge-guided general multimodal foundational model for SAR, along with reusable data and evaluation baselines. Specifically: (1) FUSAR-GEOVL-1M (the first large-scale SAR dataset with complete geographic projection attributes) was constructed, covering multiple satellite platforms, 120,000 images, and 135 cities; (2) Aligned structured text was generated through hierarchical cognitive thought chains, accurately encoding more than 1 million multidimensional semantic information from geomorphological environment and regional attributes to spatial relationships; (3) A self-consistent iterative optimization mechanism was designed to guide cross-modal learning with this knowledge information consistent with human cognition and physical laws in a self-supervised closed loop consisting of contrast, matching, and reconstruction; (4) A unified evaluation benchmark was established in 11 typical downstream tasks in the two major categories of vision and language, and compared with 15 mainstream foundation models. Experiments show that FUSAR-KLIP exhibits optimal performance, paving a new path for building remote sensing intelligent systems that are more in line with human cognitive logic. The dataset and model are publicly available at: https://github.com/yangyifremad/FUSAR-KLIP.
@@ 1 Introduction
In recent years, foundational models built around the Transformer have established a dominant position in the field of computer vision, demonstrating outstanding general representation capabilities through self-supervised paradigms such as masked image modeling or contrastive learning[10876763],[10490262]. Inspired by this, a number of foundational models targeting the characteristics of Earth observation have rapidly emerged in the field of remote sensing. To address the need for high-precision sensing, RingMo[diao2025ringmo] and SpectralGPT[10490262] have achieved deep adaptation to small targets and multispectral features through customized masked modeling and 3D spectral reconstruction. To address the challenges of wide-area generation and data sparsity, MetaEarth[yu2024metaearth] and AlphaEarth[brown2025alphaearth] have introduced generative diffusion and spatiotemporal embedding field techniques, successfully extending the model’s capabilities from passive interpretation to active simulation and multi-source assimilation, forming a diversified development pattern centered around the core needs of Earth observation[10607962],[zhou2024vlgfm],[10770756],[soni2025earthdial],[liuRandomFeatures].
While the aforementioned work has achieved record-breaking performance metrics on specific tasks, its theoretical foundation still faces significant challenges. As Muttenthaler et al. pointed out, a fundamental misalignment exists between intelligent model representations and human cognition: existing deep neural networks, although excelling in local feature matching, often fail to capture hierarchical knowledge structures like humans[muttenthaler2025aligning]. This means that current models lack the deep cognitive abilities that align with human logic, resulting in significantly limited robustness and generalization when faced with out-of-distribution data.
@@
This cognitive misalignment is particularly acute in the field of Earth observation. Because it hasn’t fundamentally broken free from the established framework of ”general vision,” existing remote sensing models tend to treat images as high-resolution natural pictures for visual understanding, neglecting the unique characteristics of remote sensing interpretation logic: remote sensing images are not merely collections of pixels, but deeply coupled entities of geomorphic features, regional attributes, and spatial structures. Lacking the ability to model hierarchical geographic information, these models often remain at a superficial level of visual perception, failing to possess deep geoscientific cognitive capabilities.
This cognitive disparity has further evolved into a dual barrier of modal heterogeneity and mechanism intricacy in the field of synthetic aperture radar (SAR)[10965804]. Although SAR possesses irreplaceable all-weather strategic value, its imaging, based on microwave coherent mechanisms, is heavily affected by speckle noise and geometric distortion[10638186]. When existing foundation models are migrated to the SAR domain, they struggle to understand the geometric structure and physical properties of ground objects through non-intuitive scattering phenomena, leading to a sharp decline in generalization performance[zhou2025fifty]. Therefore, constructing a multimodal foundation model that adapts to SAR imaging characteristics and possesses geographic cognitive capabilities has become a critical scientific imperative in the current field of remote sensing intelligence.
Realizing this vision is no easy task. Due to limitations in data privacy, domain specialization, and feature complexity, this research direction still faces many challenges, specifically:
@@
1) Geospatial Priors Remain Underutilized: Geographic information, a core element in Earth sciences, is crucial for remote sensing. However, most current SAR interpretation studies remain at the visual level, neglecting geographic attributes. This stems from the extension of natural image paradigms—focused mainly on texture and structure[Liu2016MRELBP]—as well as the closed nature of SAR data sources, with many studies still relying on outdated datasets like MSTAR from the 1990s[AFRL_MSTAR]. As shown in Table I, our review of 16 mainstream SAR datasets across tasks such as detection, segmentation, and recognition reveals a widespread omission of geographic metadata. This absence hinders regional analysis, terrain understanding, and spatial reasoning, ultimately limiting progress toward cognitive-level SAR interpretation.
@@
2) Knowledge Bottleneck: Existing remote sensing VLM studies face a significant knowledge bottleneck in text construction. On the one hand, approaches that convert detection, recognition, or segmentation annotations into templated text [10504785],[wei2025sarlang] generate descriptions that are semantically sparse and structurally repetitive, covering only shallow attributes (e.g., category and location) while neglecting complex elements such as landform, regional function, or spatial structure. This low-information text causes different images to collapse into similar semantic spaces, weakening alignment quality and fine-grained discrimination, as shown in Fig. 2(a). On the other hand, automatic text acquisition strategies, which are effective in natural image or optical remote sensing domains (e.g., web crawling, video subtitles, or AI-generated captions), fail in the SAR domain due to speckle noise and the need for expert knowledge. The resulting AI-generated text is often semantically vague or distorted Fig. 2(b). Together, these issues highlight that SAR multimodal foundational model suffers from a dual bottleneck: text derived from manual annotations lacks semantic depth, and reliable large-scale automatic acquisition remains infeasible, severely restricting scalability in open-world applications.
@@
FUSAR-MAP[shi2021object] 2021 45 C Single 1 1,024 610 TIF Segmentation
SSDD[XU2025440] 2021 633 C,X Multi $\sim1$5 500 1,160 JPG Detection
@@
FUSAR-SHIP[hou2020fusar] 2022 271 C Single 1 512 16,144 TIFF Detection
@@
SAMPLE[lewis2019sar] 2019 14 X Single 0.3 128 2,732 PNG Classification
@@
PolarSAR-Seg[wang2022air] 2022 36 C Multi 8 512 500 TIFF Segmentation
@@
In terms of text data acquisition, this paper introduces a hierarchical cognitive chain-of-thought (HCoT) instruction to simulate the human interpretation process of SAR imagery. HCoT guides GPT-4.1 [OpenAIChatGPT] to progressively incorporate multi-dimensional knowledge—such as geographic context, regional priors, SAR imaging principles, and target scale perception—enabling semantic reasoning from global to local, and from general to domain-specific knowledge. Furthermore, the constructed multi-scale image semantics driving mechanism guides the model to generate semantic expressions that connect the upper and lower scales on large, medium, and small scale images, and establish cognitive coherence. Under HCoT guidance, the language model significantly enhances its SAR interpretation ability, generating more informative and coherent textual descriptions to support multimodal representation learning.
In terms of multimodal alignment modeling, this paper proposes FUSAR-KLIP, a dual-tower multimodal framework employing Vision Transformer (ViT)[dosovitskiy2021an] and BERT[devlin2019bert] as visual and language encoders. The model jointly optimizes image-text contrastive loss (ITC), image-text matching loss (ITM), and masked language modeling loss (MLM) to construct a unified cross-modal embedding space bridging low-level perception and high-level semantics. To mitigate semantic deviations in HCoT-guided generated text, we introduce a self-consistent iterative optimization (SCIO) module that enhances alignment accuracy and stability through a closed-loop self-supervised strategy of screening, filtering, and reconstruction.
@@
  • FUSAR-KLIP (Knowledge-guided Language-Image Pre-training): The first knowledge-guided visual language foundation model for SAR was constructed, combining contrast, matching, and reconstruction multi-task learning to establish a cross-modal representation space for vision-language collaboration. The SCIO optimization module is introduced to dynamically enhance text accuracy and improve cross-modal alignment quality through the ”screen–filter–reconstruct” self-supervised closed-loop optimization mechanism.
  • Leading Multi-Task Generalization Capability: In various typical downstream tasks, such as target classification, detection, land feature segmentation, image captioning, image-text retrieval, and visual question answering, FUSAR-KLIP demonstrates superior semantic understanding and cross-task generalization performance compared to existing remote sensing multimodal models.
The remainder of this paper is organized as follows: Section 2 introduces the research progress in related fields; Section 3 details the construction process of the FUSAR-GEOVL-1M dataset and the design of the multimodal self-supervised model; Section 4 presents experimental verification and analysis across multiple remote sensing downstream tasks; Section 5 provides a summary of the paper and outlines potential future research directions.
@@ 2.1 Paradigm Evolution of Remote Sensing Interpretation
In the supervised learning stage, large-scale manual annotations enabled significant progress in recognition, detection, and segmentation tasks, but the prohibitive cost and limited semantic coverage of annotations severely constrained generalization across diverse scenes[shi2021object].
Unimodal self-supervised learning (SSL) alleviated this bottleneck by mining latent structures from unlabeled data, producing robust representations with minimal fine-tuning and laying the foundation for domain-specific models[10713915]. The RingMo series pioneered the introduction of mask learning into remote sensing, optimizing feature extraction of dense small targets in remote sensing images through targeted masking strategies[diao2025ringmo]. RingMo-Aerial further introduced a frequency-enhanced attention mechanism and affine transformation contrastive learning, solving the challenges of tilted viewpoints and multi-scale occlusion in aerial remote sensing. Addressing the 3D characteristics of multispectral data, SpectralGPT innovatively employs 3D masking and multi-target reconstruction strategies, overcoming the limitation of processing only RGB data and achieving effective capture of spectral sequence information[10490262]. In the generative direction, MetaEarth, through a self-cascaded framework and noise sampling strategy, achieved global-scale, unbounded resolution remote sensing image generation, driving the leap from perception to simulation[yu2024metaearth]. RS-vHeat takes a different approach, using a heat conduction physics model to simulate feature diffusion, significantly improving computational efficiency while maintaining the global receptive field[hu2025rs]. Furthermore, AlphaEarth proposed the concept of Embedding Fields, addressing the label sparsity problem by assimilating multi-source spatiotemporal data[brown2025alphaearth]. Nevertheless, unimodal SSL remains largely restricted to low-level perceptual cues such as texture and geometry, offering limited capacity for higher-order semantic reasoning[10915556],[10955443].
Multimodal SSL further introduces language to build cross-modal embedding spaces through large-scale image–text alignment, enabling richer semantics and task generalization. This paradigm has driven breakthroughs in natural imagery and optical remote sensing. However, SAR has received little attention within this framework: most existing models rely on RGB-based assumptions and fail to capture SAR’s unique imaging mechanisms. As a result, researchers have only begun to explore unimodal SSL in SAR as a transitional step toward multimodal expansion.
@@ 2.2 Research on Unimodal Self-supervised Foundational Models in SAR Images
SARATR-X uses a two-step SSL method with multi-scale gradient features to establish a high-performance SAR image target recognition foundational model[10856784]. Yang et al. proposed the SARDet-CL, combining feature enhancement with a SSL method constrained by imaging mechanisms, and achieved advanced performance in downstream detection tasks[10994283]. Li et al. proposed SAR-JEPA, which constructed a self-supervised pre-training foundational model for SAR target recognition tasks from the perspective of overcoming speckle noise[LI2024326]. Ren et al. introduced multi-image factor SSL to promote directional feature learning and obtain generalized features, enhancing the performance in terrain classification tasks[10496083]. Wang et al. constructed a SAR multi-task foundation model based on cross-domain continuous pre-training[10726860]. Pei et al. proposed a two-stage SAR image pre-training method based on SSL to improve the accuracy of target classification[10269665]. MSFA proposed a multi-stage filter augmentation pre-training framework for large-scale RGB and SAR data, which performed well when transferred to detection tasks[li2024sardet100k].
@@ 2.3 Research on Multimodal VLMs in Remote Sensing Field
Existing work primarily focuses on optical imagery. For instance, RemoteCLIP is based on the CLIP framework by converting class labels in public remote sensing datasets into templated text, establishing a foundational model that achieves leading performance across many typical tasks[10504785]. SkyEyeGPT developed a multimodal remote sensing command dataset and designed a two-stage tuning strategy to enhance conversational capabilities[zhan2025skyeyegpt]. VHM generated the HqDC dataset based on the large language model Gemini, easing the issue of model hallucination[pang2025vhm]. SkyScript utilized structured geographic information from OpenStreetMap to generate semantic text for optical images, improving the quality of cross-modal alignment[wang2024skyscript]. GeoChat expanded existing remote sensing image-text pairs to build a multi-round conversation dataset with command-following capabilities[kuckreja2024geochat]. BITA introduced a lightweight Fourier Transformer structure to enhance image-text interaction[10415446], while BAN developed a foundational model for remote sensing change detection tasks, leading to improved task performance[10438490]. Multimodal modeling for SAR images is still in its infancy.
In contrast, multimodal foundation model construction for SAR is still in its infancy: early attempts such as SARLANG[wei2025sarlang] and SARCLIP[jiang2026sarclip], which generates text from detection annotations, and SSL-LIP[10642118], which employs dual-stage self-supervised training with limited labels, have demonstrated feasibility but remain constrained by small-scale, template-based text construction and sparse semantics. Therefore, there is an urgent need to construct a multimodal foundation model adapted to the imaging characteristics of SAR images.
@@ 3.1 FUSAR-GEOVL-1M Image Data
The dataset covers 135 representative urban areas, as shown in Fig. 5. The scene types include typical remote sensing semantic categories such as airports, ports, urban areas, water bodies, industrial parks, and road networks. The dataset construction process consists of the following key steps:
@@
As shown in Fig. 4, the FUSAR-GEOVL-1M dataset presents a clear distribution of data, sample proportions, and representative examples. It is noteworthy that this dataset is built upon several task-oriented SAR datasets previously proposed by our team—namely, FUSAR-MAP[shi2021object], FUSAR-SHIP[hou2020fusar], and FUSAR-AIR[Qian2024Thesis]—which have undergone large-scale expansion and data restructuring. These enhancements have significantly improved the dataset’s coverage, semantic density, and adaptability to downstream tasks. The aforementioned datasets have been downloaded over 6,000 times and are widely cited and applied in the field of SAR interpretation. FUSAR-GEOVL-1M represents the systematic accumulation of our team’s long-term research in SAR image understanding and provides a strong data foundation for multimodal modeling in the SAR domain.